the perimeter of a quadrilateral is 29 inches. The longest side is twice as long as the shortest side. The other two sides are equally long and are 2 inches longer than the shortest side. Find the length of all four sides.
p.s I tried using a table to organize the variables but I can't seem to solve it.

Answer by stanbon(75887)

the perimeter of a quadrilateral is 29 inches. The longest side is twice as long as the shortest side. The other two sides are equally long and are 2 inches longer than the shortest side. Find the length of all four sides.
--------
Draw an irregular quadrilateral.
Let the shortest side be "x"
Let the longest side be "2x"
The two other sides are each "x+2"
EQUATION:
x+2x+x+2+x+2=29
5x=25
x=5 inches
2x=10 inches
x+2=7 inches
x+2=7 inches
They add up to 29 inches, as required.
Cheers,
Stan H.